Sheen Co. manufacturers laser printers. It has outlined the following overhead cost drivers: Overhead Costs Pool Quality contWhat is the total overhead cost per unit of the laser printers order using activity-based costing? (Rounded to the nearest ce51.47 is the 5th option
Answer #1

calculation of overhead cost per driver :

overhead cost pool cost driver overhead cost budget level of cost driver overhead cost per driver
quality control no. of inspection $90000 1500 $90000/1500 = $60
machine operations machine hours $156000 1500 $156000/1500 = $104
material holdings no. of batches $1200 40 $1200/40 = $30
misc. overhead cost direct labor hours $64000 4000 $64000/4000 =$16

calculation of total overhead cost of 1000 laser printers :

cost drivers production requirement overheadd cost per driver total overhead cost
no. of inspection 200 $60 200 x $60 = $12000
machine hours 255 $104 255 x $104 = 26520
no. of batches 5 $30 5 x $30 =$150
direct labor hours 800 $16 800 x $16 = $12800
total overhead costs $51470

total overhead cost per unit = total overhead cost/no. of units = $51470/100 laser printers = $51.47 per laser printer

as per my calculations correct answer must be $51.47 which is not available in options.

please make sure given options are complete or missing or correct.

thank you

please give your feed back and rating